You are selling tickets for a musical at your local community college. Student tickets cost $5 and general admision tickets cost $8. If you sell 500 tickets and collect $3475, how many student tickets and how many general admission. Thank you for your help and time .
Answer by Nate(3500)

student tickets = s
adult tickets = a
a + s = 500 or a = 500 - s
5s + 8a = 3475
5s + 8(500 - s) = 3475
5s + 4000 - 8s = 3475
-3s = -525
s = 175
a = 500 - s
a = 500 - 175 = 325